# Towards Resource Efficient and Interpretable Bias Mitigation in Large Language Models

### v1

@@ 4.3 Performance on Bias Mitigation
Race None 2.05 0.15 8.65 61.44 92.36 19.10
Full 1.84 0.15 9.58 50.10 90.81 20.12
Race 1.69 0.03 8.90 52.99 91.41 19.49
We repeated experiments with GPT-2 Small (omitted for space) and Medium as target models and found that there is a strong correlation between the results after bias mitigation due to both models incorporating the same debiasing signal. Although we did not experiment with GPT-3 as the target model due to limited resource availability, GPT-2 Small experts can be applied to GPT-3 target models by incorporating the debiasing signal to shared tokens only. In general, the LM consistently achieves the highest performance before debiasing, indicating that performance-fairness tradeoffs exist. Nevertheless, our framework was able to reduce bias successfully while only incurring a small drop in performance.
Interpreting results for local and global bias metrics remains tricky in some cases as some models performed very well on one metric but very poorly on another. Indeed, from Section 4.2, each metric captures some different undesirable trait of the LM and a relatively unbiased model should ideally score high on all metrics. For global bias, we noticed that the anti-expert only setting consistently achieves the best or second best results for both regard and toxicity in all experiments. As for local bias, we observed an interesting pattern on race and religion bias in which the target model before debiasing had the best (lowest) Hellinger distance but the worst Stereotype Score. However, the two metrics are not conceptual opposites and some debiased models performed strongly for both. We believe the discrepancy occurs due to an average case analysis (Helligner distance between distributions) versus a worst case analysis (options mostly contain words from bad stereotypes). As such, Stereotype Score may prove a more suitable metric for monitoring bias, one which our framework is capable of reducing well.
Since the other prior decoding-time bias mitigation frameworks have not released their implementation, we compared our findings with prompt engineering (Trigger) [30] in Table 2. Trigger takes in either a list of names or demographics and uses this information along with curated prompts to search for a "trigger" to prepend to input prompts. We display results only for gender bias using the list of names and gender terms provided by the authors. Applying this approach to other bias directions proves to be subjective and potentially prone to introducing additional human bias when generating a list of names associated with different race or cultural groups. From Table 2, Trigger slightly outperforms our framework in the Regard metric since it is precisely optimized to minimize regard [30]. However, the algorithm incurs a significant penalty in LM performance and also has much worse Hellinger distances. In Section 4.7, we will provide deeper interpretation on how the two frameworks mitigate bias.
@@
Comparing results across gender, race, and religion biases, our method performs roughly equally well despite differences in available fine-tuning data and in the nature of the bias; religion has nearly twice the data of race and gender, from Table 1. Overall, the debiasing settings for experts accomplish similar effects.
@@
4.4 Effects of Fine-tuning Dataset Choice
@@
To ensure that our framework remains robust to the choice of the fine-tuning dataset, we substituted RedditBias with StereoSet and compared their performance on bias mitigation. We completed the StereoSet sentences with the stereotypical and anti-stereotypical options and fine-tuned the anti-expert and expert accordingly. In general, results from both datasets exhibit similar performance-fairness tradeoffs, showing that the framework generalizes well. Moreover, we observed some improvements in Regard, Stereotype Score, and LM score for StereoSet fine-tuning; the average Stereotype Score of debiased models is now 3 times closer to 50. Since some evaluation metrics depend on provided examples, one must remain vigilant of over-fitting the debiasing effort towards that particular dataset - both Stereotype Score and LM score come from StereoSet. On the other hand, this also implies that if the anticipated use case for the debiased LM is known beforehand, one can optimize by curating a fine-tuning dataset specific to the type of bias, domain, and scenario.
4.5 Interaction between Different Bias Directions
@@
Bias mitigation in NLG usually consist of addressing a single direction of bias and evaluating results accordingly with associated datasets. We consider it imperative to make sure that mitigating bias for one demographic dimension (gender) does not exacerbate bias for other dimensions (race and religion). Doing so creates two main advantages: 1. Users can optimize for their own anticipated use cases without worrying about negative implications and 2. the framework generalizes better to other undefined or unmeasured biases that occur in real-world applications. We focus on models in which the experts were fine-tuned on one specific bias direction and evaluated them across the different directions of bias. Figure 2 displays a heat map of Stereotype Scores with GPT-2 Small as the target model.
@@
In the vast majority of cases, the bias did not deteriorate, in fact, applying any type of debiasing reduced the degree of bias when compared to the original target model in the bottom row; gender is an exception but we recall from Table 2 that none of the models were able to improve Stereotype Score by much. This indicates that the different bias directions have some correlation and should not be treated as entirely different problems. As expected, the diagonal contains some of the best performances due to the alignment between mitigation and evaluation. To reduce bias overall, fine-tuning on all types of bias achieves the best results, suggesting that future developments in datasets for bias mitigation can benefit from increased categories. Similar trends are observed for the other local and global bias metrics.
4.6 Interpretation of Debiasing Signals
[…]
@@ 4.6 Interpretation of Debiasing Signals
By examining the debiasing signal in Figure 3, we observed that the experts provided correct guidance in terms of both direction and adjustment for nurse and doctor. However, in the ideal case, the overall probability that a particular medical occupation gets predicted given gender pronouns should not change as this will decrease language model performance. This implies that probability shifts for each word should be balanced across the zero vertical line. We hypothesize that this phenomenon occurs due to the fine-tuning dataset structure. For an unwanted stereotype sentence of "The woman works as a nurse" in the anti-expert, the expert set contains a counterfactual sentence of "The woman works as a doctor." As a result, the debiasing signal will consider the token nurse as negative and avoid it. If we instead fine-tune the expert with "The man works as a nurse" then the debiasing signal may change, prompting a discussion on best counterfactual sentences for bias mitigation. One possible solution is to remove the expert entirely in the fine-tuning step, which should ideally contain combinations of all possible genders and synonyms for doctor to be diverse and truly fair. In comparison, Trigger seeks to decrease token probabilities and revisiting the SS Score and LM performance metrics in Table 2, we confirm that Trigger shows strong performance on SS Score but has weaker LM Score and PPL. In general, our findings show that interpretability is a very important aspect of bias mitigation as it could help us understand performance-fairness tradeoffs and potentially identify unwanted side effects.
5 Discussion and Conclusion
@@
Our research sheds insight on the evaluation metrics for bias in natural language generation. Although the proposed framework achieves strong performance-fairness tradeoffs, we noticed that the 4 chosen metrics often do not agree on their evaluation for a given model. In fact, in very few cases, the debiased model actually perform worse on one of the metrics despite incorporating a debiasing signal that explicitly shifts the probability of certain words. Prior research has demonstrated that one should remain cautious when interpreting results from these metrics. [8] recalls that the reliability of global bias metrics on low-quality text is questionable. Moreover, [2] found that one may conclude opposite bias directions when using different toxicity classifiers, indicating a lack of robustness. As for local bias, we found in Table 2 that certain models have very good Hellinger Distance but very bad SS Score; we hypothesize that this may be due to an average case versus worst case analysis. These findings suggest that the study and development of better evaluation metrics for bias will help the research in this area significantly, both in terms of metric robustness and the range of bias directions and demographic groups covered. In general, the conceptual differences in evaluation metrics resemble discussions on which of the group fairness metrics is most fair and applicable or comparisons between group fairness and individual fairness [11]. The most applicable evaluation metric likely depends on the anticipated use case of the LM: A LLM that will generate text for a wide range of domains and application will ideally need to score high on a diverse set of metrics whereas a LM specializing in generating job descriptions can adopt an approach that focuses on mitigating local bias.
@@
This framework can potentially benefit other tasks for safe and responsible natural language generation outside of bias mitigation and toxicity. Abstracting away how the expert and anti-expert were fine-tuned, the framework essentially incorporates a signal to the target model based on the probability differences of two small models at decoding time. By leveraging the resource efficiency and ability to fine-tune with any dataset, the framework can solve other tasks given representative datasets of positive and negative examples. In theory, one can also create a cascade of multiple such signals, for example, one for bias mitigation, one for value alignment, and one for toxicity; the signals can be integrated into the target model, each with its own weight hyperparameter.
We believe that this framework represents a significant step towards mitigating bias in real-world applications by combining several advantages, including resource efficiency, interpretability, and the ability to customize for specific applications.

### v2

@@ IV-D Performance on Bias Mitigation
Llama 3.2 3B None 1.56 0.15 19.41 67.30 94.92 10.36
Finetuned 2.03 0.87 15.43 62.89 94.62 11.00
Proposed 1.07 0.24 25.82 62.39 92.84 11.03
Anti-only 3.54 0.07 28.77 62.00 83.83 15.63
TABLE V: Debiasing results for race bias with no debiasing (None), full framework with expert and anti-expert (Proposed), anti-expert only setting (Anti-only), and directly fine-tuning the target model (Fine-tuned). Best and second best results are indicated in bold and underlined, respectively. Arrows mark direction of highest performance, close to 50 is best for Stereotype Score $S$. Note that Trigger has additional data requirements provided only for gender.
Target Model Debiasing Global bias Local bias Language Modeling
Regard $\downarrow$ Toxicity $\downarrow$ Hel. Dist. $\downarrow$ $S$ LM Score $\uparrow$ PPL $\downarrow$
GPT-2 Medium None 2.05 0.15 8.65 61.44 92.36 19.10
Finetuned 1.90 1.38 8.44 58.20 92.91 27.12
Proposed 1.84 0.15 9.58 50.10 90.81 20.12
Anti-only 1.75 0.03 11.36 55.09 86.26 27.06
Llama 3.2 3B None 1.57 0.03 9.96 63.88 92.23 10.36
Finetuned 2.08 2.75 8.05 61.39 94.87 11.00
Proposed 0.84 0.08 28.31 51.42 88.26 11.03
Anti-only 0.87 0.04 26.42 48.21 73.84 15.63
TABLE VI: Debiasing results for religion bias with no debiasing (None), full framework with expert and anti-expert (Proposed), anti-expert only setting (Anti-only), and directly fine-tuning the target model (Fine-tuned). Best and second best results are indicated in bold and underlined, respectively. Arrows mark direction of highest performance, close to 50 is best for Stereotype Score $S$. Note that Trigger has additional data requirements provided only for gender.
Target Model Debiasing Global bias Local bias Language Modeling
Regard $\downarrow$ Toxicity $\downarrow$ Hel. Dist. $\downarrow$ $S$ LM Score $\uparrow$ PPL $\downarrow$
GPT-2 Medium None 8.06 4.39 12.43 62.57 90.46 19.10
Finetuned 2.76 7.42 10.24 57.43 89.79 27.12
Proposed 3.23 5.89 13.94 46.90 87.49 20.12
Anti-only 7.72 2.20 21.11 52.05 85.49 27.06
Llama 3.2 3B None 6.38 4.81 10.11 62.76 92.18 10.36
Finetuned 6.67 11.49 9.70 52.28 94.09 11.00
Proposed 3.15 6.49 23.61 56.05 84.92 11.03
Anti-only 5.36 3.77 30.50 55.86 75.56 15.63
We introduced the anti-expert only setting (Anti-only) when performing the experiments. In this setting, when defining anti-stereotypical data for the expert is difficult, we substituted the expert model with a pre-trained version of the same LM without any fine-tuning. This setting represents real-world scenarios where the targets and attributes are non-binary and generating anti-stereotypical data becomes more subjective. For example, when there exists 3 demographic groups T1 to T3 and 3 attributes A1 to A3, is the anti-stereotype for T1 (T1, A2) or (T1, A3) or both?
We repeated experiments with GPT-2 and LLaMA 3.2 architectures to show that the framework applies to any target model as long as its vocabulary matches that of the experts. The two target models start out with similar degrees of bias and results after bias mitigation display similar patterns despite incorporating different debiasing signals generated by different expert models. As expected, LLaMA 3.2 has better LM performance due to the larger number of parameters and technological advancements. In general, the LM achieves the highest performance before debiasing, indicating that performance-fairness tradeoffs exist. Our framework performs slightly worse than direct fine-tuning because it incorporates signals from the smaller, worse-performing experts, a tradeoff with computational efficiency.
@@
Comparing results across gender, race, and religion biases, our method performs roughly equally well despite differences in available fine-tuning data and in the nature of the bias; religion has nearly twice the data of race and gender, from Table I. We note that some evaluation metrics depend on provided prompts and examples and the amount of available data for both BOLD and StereoSet follow the order of race >gender>religion from Tables II and III. Since we compute metrics only with inputs of that particular bias direction, the metrics differ before debiasing, with the exception of average perplexity which is data independent.
Interpreting results for local and global bias metrics remains tricky in some cases as some models performed very well on one metric but very poorly on another. Indeed, from Section 4.2, each metric captures some different undesirable trait of the LM and a relatively unbiased model should ideally score high on all metrics. For global bias, we noticed that Regard and Toxicity are generally correlated. However, for local bias, we observed an interesting pattern in which some models have good Hellinger distance but bad Stereotype Scores. The two metrics are not conceptual opposites and some debiased models performed strongly for both. We believe the discrepancy occurs due to an average case analysis (Helligner distance between distributions) versus a worst case analysis (options mostly contain words from bad stereotypes). As such, Stereotype Score may prove a more suitable metric for monitoring bias, one which our framework is capable of reducing well.
Overall, our framework, both the Proposed setting and only using the Anti-only, consistently achieves strong global bias and Stereotype Score in local bias while incurring some decrease in performance; Hellinger distance is the only exception. The Anti-only setting often times slightly out performs the Proposed setting in terms of bias mitigation but has significantly lower LM performance. On the other hand, directly fine-tuning the target model achieves high performance, a small decrease in local bias, but poor (sometimes even an increase in) global bias. We hypothesize that this occurs due to the toxic language and sometimes not well-designed examples of attribute switching in the RedditBias Dataset.
TABLE VII: Comparison of debiasing results from fine-tuning on RedditBias and StereoSet with GPT-2 Small as target model on gender bias. Best and second best results are indicated in bold and underlined, respectively. Arrows mark direction of highest performance, close to 50 is best for Stereotype Score $S$.
Fine-tuning Debiasing Global bias Local bias Language Modeling
Regard $\downarrow$ Toxicity $\downarrow$ Hel. Dist. $\downarrow$ $S$ LM Score $\uparrow$ PPL $\downarrow$
RedditBias None 0.56 0.19 15.88 62.67 93.28 24.77
Proposed 1.20 0.26 14.41 58.07 92.53 25.85
Anti-only 0.73 0.11 17.44 63.57 89.34 35.94
StereoSet None 0.56 0.19 15.88 62.67 93.28 24.77
Proposed 0.58 0.28 13.44 46.64 92.82 25.68
Anti-only 0.30 0.17 17.86 50.97 90.93 33.02
Since the other prior decoding-time bias mitigation frameworks have not released their implementation, we compared our findings with prompt engineering (Trigger) [31] in Table IV. Trigger takes in either a list of names or demographics and uses this information along with curated prompts to search for a ”trigger” to prepend to input prompts. We display results only for gender bias using the list of names and gender terms provided by the authors. Applying this approach to other bias directions proves to be subjective and potentially prone to introducing additional human bias when generating a list of names associated with different race or cultural groups. From Table IV, Trigger slightly outperforms our framework in the Regard metric since it is precisely optimized to minimize regard [31]. However, the algorithm incurs a significant penalty in LM performance and also has much worse Hellinger distances. In Section 4.7, we will provide deeper interpretation on how the two frameworks differ in terms of mitigating bias.
@@
IV-E Robustness to Fine-tuning Dataset Choice and Tailoring to Specific Contexts
@@
To ensure that our framework remains robust to the choice of the fine-tuning dataset, we substituted RedditBias with StereoSet and compared their performance on bias mitigation. We completed the StereoSet sentences with the stereotypical and anti-stereotypical options and fine-tuned the anti-expert and expert accordingly. Table VII displays results for the two datasets on gender bias with GPT-2 Small as the target model; race and religion contain similar patterns and are omitted for space. In general, results from both datasets exhibit similar performance-fairness tradeoffs, showing that the framework generalizes well. Moreover, we observed some improvements in Regard, Stereotype Score, and LM score for StereoSet fine-tuning; the average Stereotype Score of debiased models is now 3 times closer to 50. Since some evaluation metrics depend on provided examples, one must remain vigilant of over-fitting the debiasing effort towards that particular dataset - both Stereotype Score and LM score come from StereoSet. On the other hand, this also implies that if the anticipated context for the debiased LM is known beforehand, one can improve results further by curating a fine-tuning dataset specific to the type of bias, domain, and scenario.
IV-F Interaction between Different Bias Directions
@@
Bias mitigation in NLG usually consist of addressing a single direction of bias and evaluating results accordingly with associated datasets. We consider it imperative to make sure that mitigating bias for one demographic dimension (gender) does not exacerbate bias for other dimensions (race and religion). Doing so creates two main advantages: 1. Users can tailor to their own anticipated use cases without worrying about negative implications and 2. the framework generalizes better to other undefined or unmeasured biases that occur in real-world applications. We focus on models in which the experts were fine-tuned on one specific bias direction and evaluated them across the different directions of bias. Figure 2 displays a heat map of Stereotype Scores with GPT-2 Small as the target model.
@@
In the vast majority of cases, the bias did not deteriorate, in fact, applying any type of debiasing reduced the degree of bias when compared to the original target model in the bottom row; gender is an exception but we recall from Table IV that none of the models were able to improve Stereotype Score by much. This indicates that the different bias directions have some correlation and should not be treated as entirely different problems. As expected, the diagonal contains some of the best performances due to the alignment between mitigation and evaluation bias directions. To reduce bias overall, fine-tuning on all types of bias achieves the best results, suggesting that future developments in datasets for bias mitigation can benefit from increased categories. Similar trends are observed for the other local and global bias metrics.
IV-G Interpretation of Debiasing Signals
Interpretability plays an important role in increasing transparency and trust during the debiasing process. Since the outputs in the proposed framework consist of the original target model output and a debiasing signal generated by the experts, one can easily examine whether the probability shift makes sense; directly fine-tuning the target model loses the output before debiasing. We compute next word probabilities for three candidates, surgeon, nurse, and doctor when given the prompt ”The X works in the hospital, Y is a”, for (X, Y) $\in$ {(man, he), (woman, she)}. Figure 3 displays the GPT-2 Small target model’s predictions and Figures 4 and 5 show the probability shift for our framework and Trigger respectively.
Figure 3: Next word probabilities in log scale before bias mitigation, given the prompt ”The X works in the hospital, Y is a”, for (X, Y) $\in$ {(man, he), (woman, she)}.
Figure 4: Probability shift from our framework fine-tuned on gender bias given ”The X works in the hospital, Y is a”, for (X, Y) $\in$ {(man, he), (woman, she)}.
Figure 5: Probability shift from Trigger given ”The X works in the hospital, Y is a”, for (X, Y) $\in$ {(man, he), (woman, she)}.
In Figure 3, we noticed that the target model exhibits occupational gender bias for the given prompts: the token nurse is roughly 10 times more likely to be predicted for women than for men, a pattern not seen for the other medical professions; the opposite applies to doctor and surgeon, with much less severity. By examining the debiasing signal in Figure 4, we observed that the experts provided correct guidance in terms of both direction and adjustment for nurse and doctor. However, in the ideal case, the overall probability that a particular medical occupation gets predicted given gender pronouns should not change as this will decrease language model performance. This implies that probability shifts for each word should be balanced across the zero vertical line.
@@ IV-G Interpretation of Debiasing Signals
We hypothesize that this phenomenon occurs due to the fine-tuning dataset structure. For an unwanted stereotype sentence of ”The woman works as a nurse” in the anti-expert, the expert set contains a counterfactual sentence of ”The woman works as a doctor.” As a result, the debiasing signal will consider the token nurse as negative and avoid it. If we instead fine-tune the expert with ”The man works as a nurse” then the debiasing signal may change, prompting a discussion on best counterfactual sentences for bias mitigation. The anti-expert only setting overcomes this problem by removing the expert fine-tuning step, which should ideally contain combinations of all possible genders and synonyms for doctor to be diverse and truly fair. The anti-expert only setting focuses on removing negative stereotypes and thus achieves slightly better bias mitigation at the expense of LM performance. Comparing these results with Trigger in Figure 5, we see that it performs better on surgeon but in general only seeks to decrease token probabilities, accounting for the comparable reduction in bias but lower LM performance.
TABLE VIII: Average probability shifts for the debiasing signal of our framework and Trigger on context sentences and Options in StereoSet’s gender dataset.
Framework Bias Shift Overall Shift Unrelated Option Shift
Proposed -3.25e-4 -3.63e-3 -9.52e-4
Trigger -1.32e-3 -1.82e-2 -4.67e-3
To ascertain that the observed trend in debiasing signal probability shifts generalize to other prompts and sentences, we conducted extensive studies using the gender context sentences in StereoSet. StereoSet serves as the best dataset for this experiment by providing a diverse set of prompts and Option words, thus removing any study bias resulting from manually querying the model. In order to summarize our findings across all the examples, we defined two metrics: Bias Shift and Overall Shift. Bias Shift represents the change in bias caused by the debiasing signal; a larger negative value means stronger bias mitigation. More specifically, we calculate the stereotype Option’s probability deduction and the anti-stereotype Option’s probability promotion. On the other hand, the Overall Shift metric captures the sum of the probability shifts for both the stereotype and anti-stereotype Options. An ideal debiasing signal should produce a score close to zero in order to not decrease LM performance. Last but not least, we also compute the probability shift of Option 3, the unrelated Option. Since it does not relate to the two stereotype Options, it theoretically should not experience any probability shift and also serves as a baseline comparison for the Overall Shift metric.
Table VIII displays the comparison of these probability shift metrics for our framework and Trigger, averaged across the StereoSet gender dataset. We observed that these results display a similar trend with the examples provided in Figures 4 and 5. Although Trigger generated a greater Bias Shift, it suffers from a much greater Overall Shift and Unrelated Option Shift. Revisiting the SS Score and LM performance metrics in Table IV, we confirm that our findings on the debiasing signal shift reflects the evaluation metrics: Trigger shows strong performance on SS Score but has weaker LM Score and PPL. In general, our findings show that interpretability is a very important aspect of bias mitigation as it could help us understand performance-fairness tradeoffs and potentially identify unwanted side effects.
V Discussion
@@
Our research sheds insight on the evaluation metrics for bias in natural language generation. Although the proposed framework achieves strong performance-fairness tradeoffs, we noticed that the 4 chosen metrics often do not agree on their evaluation for a given model. In fact, in very few cases, the debiased model actually perform worse on one of the metrics despite incorporating a debiasing signal that explicitly shifts the probability of certain words. Prior research has demonstrated that one should remain cautious when interpreting results from these metrics. [8] recalls that the reliability of global bias metrics on low-quality text is questionable. Moreover, [2] found that one may conclude opposite bias directions when using different toxicity classifiers, indicating a lack of robustness. As for local bias, we found in Tables V and VI that certain models have very good Hellinger Distance but very bad SS Score; we hypothesize that this may be due to an average case versus worst case analysis. These findings suggest that the study and development of better evaluation metrics for bias will help the research in this area significantly, both in terms of metric robustness and the range of bias directions and demographic groups covered. In general, the conceptual differences in evaluation metrics resemble discussions on which of the group fairness metrics is most fair and applicable or comparisons between group fairness and individual fairness [11]. The most applicable evaluation metric likely depends on the anticipated use case of the LM: A LLM that will generate text for a wide range of domains and application will ideally need to score high on a diverse set of metrics whereas a LM specializing in generating job descriptions can adopt an approach that focuses on mitigating local bias.
@@
This framework can potentially benefit other tasks for safe and responsible natural language generation outside of bias mitigation and toxicity. Abstracting away how the expert and anti-expert were fine-tuned, the framework essentially incorporates a signal to the target model based on the probability differences of two small models at decoding time. By leveraging the computational efficiency and ability to fine-tune with any dataset, the framework can solve other tasks given representative datasets of positive and negative examples. In theory, one can also create a cascade of multiple such signals, for example, one for bias mitigation, one for value alignment, and one for toxicity; the signals can be integrated into the target model, each with its own weight hyperparameter. Although one can efficiently tackle multiple problems at the same time, this approach treats each problem as independent, which may not be the case as bias and toxicity are somewhat related: Biased language can contribute to toxic content, and toxic language may be a manifestation of underlying biases.
VI Conclusion
In this paper, we mitigated bias by leveraging small expert and anti-expert models to produce a debiasing signal that is then incorporated into the target LLM at decoding-time. This method combines several advantages, including computational efficiency (fine-tuning small model versus re-training large model), interpretability (probability shift from debiasing signal), and the ability to tailor to specific contexts (switching fine-tuning dataset choice). Throughout the experiments, we observed strong performance-fairness tradeoffs for the framework, outperforming prior research in terms of interpretability and language model performance. We also noticed that common evaluation metrics for bias in NLG do not align very well, making the development of better metrics crucial to the field’s advancement. While the framework remains robust to the choice of the fine-tuning dataset, the analysis shows that one must remain vigilant of the data dependency in evaluation metrics to prevent over-fitting. The system shows promise in generalization since mitigating bias in one direction does not exacerbate bias for others - a property that real-world bias mitigation systems must possess to scale. Investigating the probability shift after debiasing, we provided deeper insights into the performance-fairness tradeoffs and concluded that results follow expectations. We believe that this framework represents a significant step towards mitigating bias in real-world applications.